OpenStudy (anonymous):

Which graph best shows the function f(x) = 3(3)-x to represent the rate at which a radioactive substance decays?

OpenStudy (anonymous):

this is the equation of a line f(x)=9-x y=9-x You could find the graphic giving some valors. For example: For x=0 which is the valor of y? y=9-0=9 so , now you have one point of your line (0,9) For y=0 0=9-x x=9 Other point (9,0) 

OpenStudy (anonymous):

you guys have got to draw out your equations, or use the ^ symbol for exponents and use parentheses.

OpenStudy (anonymous):

Pretty sure that's supposed to be an exponential function not linear

OpenStudy (anonymous):

lol